# Thermally Activated Dual-Modal Adversarial Clothing against AI Surveillance Systems

### v1

@@ 1 Introduction
Figure 1: Comparisons of representative adversarial patch attacks. (a) shows that most prior adversarial patches are single-modal (RGB or infrared) and always-on in everyday settings, making them more noticeable in real-world scenarios. (b) presents that our adversarial patch attack achieves dual-modal visible-infrared deception, and support controllable activation under varying environmental conditions.
@@
Among anti-AI technologies, adversarial patch is a representative proactive countermeasures against AI surveillance [19-hingun2023reap, 20-hu2025two, 21-wei2023unified, 22-long2025cdupatch, 23-brown2017adversarial, 24-thys2019fooling, 25-wei2024revisiting, 26-liu2018dpatch, 27-long2024papmot, liu2025eva]. The patch is carefully designed visual pattern can be physically attached to clothing or belongings, causing AI models to misclassify or entirely overlook the wearer in the real world. In recent years, privacy advocates have increasingly adopted adversarial patches as a form of anti-surveillance fashion, integrating them into wearable clothing to safeguard personal data in public spaces [28-xu2020adversarial, 49-adversarialfashion, 50-capable2025clothing]. However, existing adversarial patches are often high-saturation, high-contrast textures that maximize feature perturbations. While effective against AI detectors, such conspicuous patterns easily attract human attention, making them impractical for everyday social environments. Therefore, balancing attack effectiveness with real-world usability has emerged as a key challenge in adversarial patch design.
@@ 2.2 Adversarial patch attacks
Previous digital attacks [jia2021iou, jia2022exploring, liu2025lighting, long2025robust, huang2024towards] are challenging to deploy in the physical world, as their perturbations are often imperceptible to the human eye and difficult to faithfully reproduce. In contrast, adversarial patches have emerged as a widely adopted physical attack method against AI systems [19-hingun2023reap, 20-hu2025two, 21-wei2023unified, 22-long2025cdupatch, 23-brown2017adversarial, 24-thys2019fooling, 25-wei2024revisiting, 26-liu2018dpatch, 27-long2024papmot, zhu2024infrared, zhu2025physical]. Early patch works focused on single-modal attacks. Thys et al. [24-thys2019fooling] demonstrate physically realizable adversarial patches that successfully mislead person detectors in the visible domain, while Zhu et al. [zhu2024infrared] propose infrared adversarial car stickers generated via a 3D mesh–shadow optimization to deceive thermal detectors. More recently, several works have extended patch attacks to visible–infrared dual-modal detectors [20-hu2025two, 22-long2025cdupatch, 21-wei2023unified]. For example, Long et al. [22-long2025cdupatch] propose CDUPatch, a color-driven cross-modal adversarial patch that achieves unified visible-infrared attacks by mapping RGB colors to corresponding thermal responses. Wei et al. [21-wei2023unified] propose a unified adversarial patch that simultaneously fools dual-modal detectors by optimizing patch shape with boundary-limited shape optimization, demonstrating high cross-modal physical-world attack success. However, most existing patch attacks are always-on in everyday settings, which increases their conspicuity and complicates real-world deployment. In this paper, we co-design a thermally activated patch to mislead both visible- and infrared-spectrum detectors while remaining concealed within ordinary clothing. By leveraging temperature as a control signal, our system activates both the visible and infrared adversarial patterns simultaneously, enabling controllable, on-demand dual-modal evasion.
@@ 3.2 Structure of adversarial clothing
To enable proactive privacy protection under AI surveillance, we design an adversarial clothing featuring on-demand activation and dual-modal visible-infrared camouflage. As shown in Figure 2(a), before activation the person is detected by infrared- and visible-spectrum detectors with high confidence (0.84 and 0.85, respectively); after activation, both detectors fail to identify the person (confidence << 0.4). Thermal contrast perturbs the infrared-spectrum detector, while color change deceives the visible-spectrum detector, as demonstrated in the corresponding activation video provided in Supplementary Materials. Figure 2(b) illustrates the layered structure of the clothing: (i) a thermochromic layer that switches from black to colorless when heated ($\sim$30 ºC), (ii) an adversarial patch layer carrying the algorithmically optimized RGB texture, (iii) a flexible silicone heating layer with precise temperature control, which activates the thermochromic layer and generates controlled infrared patterns, and (iv) a fabric substrate that provides thermal insulation for the wearer. Detailed descriptions of each layer are provided in the following subsections.
@@ 3.5 Adversarial patch layer
In the texture update phase (Figure 5b), we fix the optimized shape and learn a color texture that deceives the visible-spectrum detector. To account for real-world variations, we adopt the Expectation over Transformation [48-athalye2018synthesizing] (EOT) to simulate real-world transformations (e.g., rotation, brightness changes, blur, and scaling), thereby improving the patch’s attack effectiveness under diverse physical conditions. Then, the transformed patch is directly pasted onto RGB images to form visible-spectrum adversarial images. Finally, these patched images are fed into the visible detector, and the patch texture is updated via adversarial losses. We adopt a combination of total variation loss ($\mathcal{L}_{tv}$) and average precision loss ($\mathcal{L}_{ap}$) as the final adversarial objective, which is widely used in adversarial patch optimization [24-thys2019fooling, 27-long2024papmot, 28-xu2020adversarial].
@@
Figure 6: Comparison of attack success rates (ASR) for various adversarial patch methods on infrared and visible datasets. Columns from left to right: T-SEA [36-huang2023t], AdvYOLO [24-thys2019fooling], CAP [25-wei2024revisiting], FDA [37-cheng2024full], AdvTexture [30-hu2022adversarial], AdvCloak [38-wu2020making], our RGB patch, AdvIC [31-hu2024adversarial], AdvIB [32-hu2024adversarial], AIP [33-wei2023physically], HIC-IR [34-zhu2023hiding], Bulb [35-zhu2024hiding], our infrared patch. It shows that our dual-modal patch achieves the highest attack success rate compared to other methods.
@@ 4.1 Experimental Settings
Datasets and victim models. For digital attack experiments, we evaluate our adversarial patch on person detection benchmarks using the FLIR [40-flir_dataset], INRIA [41-dalal2005histograms], PennFudan [42-wang2007object], LLVIP [39-jia2021llvip] datasets. The FLIR and LLVIP dataset are used to assess attack performance under infrared spectrum detection, while the INRIA and PennFudan datasets are used for evaluation in the visible spectrum. We evaluate our attack on three mainstream detection paradigms: one-stage detectors (i.e., YOLOv3 [43-redmon2018yolov3], YOLOv5 [44-jocher2020yolov5]), two-stage detectors (i.e., Faster R-CNN [45-ren2015faster]) and transformer-based detectors (i.e., DETR [46-carion2020end]). To ensure accurate detection of both infrared- and visible-spectrum pedestrian targets, we use the officially pre-trained weights using the COCO [47-lin2014microsoft] dataset as the initial weights and then fine-tune these models on the training sets of INRIA and FLIR datasets. After fine-tuning, all detection models achieve over 90% mean Average Precision (mAP) on the their validation set.
@@
Figure 7: Real-world testing in diverse AI surveillance scenarios, including indoor room, shopping mall, and outdoor street. It proves that our adversarial achieves dual-modal evasion across both visible and infrared detectors. Refer to Supplementary Materials for video demos.
@@ 4.2 Main Results
Quantitative analysis. Figure 6 compares the performance of various adversarial patch methods on the INRIA and LLVIP Datasets. As can be seen, our dual-modal patch consistently achieves the highest attack success rate (ASR) across four mainstream detectors (i.e., Yolov3 [43-redmon2018yolov3], Yolov5 [44-jocher2020yolov5], FastRCNN [45-ren2015faster], DETR [46-carion2020end]), outperforming both visible patch methods (i.e., T-SEA [36-huang2023t], AdvYOLO [24-thys2019fooling], CAP [25-wei2024revisiting], FDA [37-cheng2024full], AdvTexture [30-hu2022adversarial], AdvCloak [38-wu2020making]) and infrared patch methods (AdvIC31, AdvIB [32-hu2024adversarial], AIP [33-wei2023physically], HIC-IR [34-zhu2023hiding], Bulb [35-zhu2024hiding]). For example, on the INRIA dataset under visible-spectrum detection, our patch attains ASRs of 54.4% (YOLOv3), 49.9% (YOLOv5), 46.8% (Faster R-CNN), and 38.4% (DETR), surpassing AdvYOLO (34.6%, 31.4%, 29.7%, and 25.2%, respectively). On the LLVIP dataset under infrared detection, our patch achieves 85.4%, 88.6%, 83.3%, and 80.2%, compared to AdvIB’s 72.7%, 75.1%, 70.3%, and 65.8%. Unlike prior square-shaped patches (e.g., CAP [25-wei2024revisiting], AdvYOLO [24-thys2019fooling]), our design employs an algorithmically optimized polygonal geometry. This irregular shape enhances patch effectiveness against both visible- and infrared-spectrum detectors, allowing superior performance across diverse datasets and complex environments.

### v2

@@ 1 Introduction
Figure 1: Comparisons of representative adversarial patch attacks. (a) shows that most prior adversarial patches are single-modal (RGB or infrared) and always-on in everyday settings, making them more noticeable in real-world scenarios. (b) presents that our adversarial patch attack achieves dual-modal visible-infrared deception, and support controllable activation in the real world.
@@
Among anti-AI technologies, adversarial patch is a representative proactive countermeasures against AI surveillance [19-hingun2023reap, 20-hu2025two, 21-wei2023unified, 22-long2025cdupatch, 23-brown2017adversarial, 24-thys2019fooling, 25-wei2024revisiting, 26-liu2018dpatch, 27-long2024papmot, liu2025eva]. The patch is carefully designed visual pattern can be physically attached to clothing or belongings, causing AI models to misclassify or entirely overlook the wearer in the real world. In recent years, privacy advocates have increasingly adopted adversarial patches as a form of anti-surveillance fashion, integrating them into wearable clothing to safeguard personal data in public spaces [28-xu2020adversarial, 49-adversarialfashion, 50-capable2025clothing, hu2023physically]. However, existing adversarial patches are often high-saturation, high-contrast textures that maximize feature perturbations. While effective against AI detectors, such conspicuous patterns easily attract human attention, making them impractical for everyday social environments. Therefore, balancing attack effectiveness with real-world usability has emerged as a key challenge in adversarial patch design.
@@ 2.2 Adversarial patch attacks
Previous digital attacks [jia2021iou, jia2022exploring, liu2025lighting, long2025robust, huang2024towards] are challenging to deploy in the physical world, as their perturbations are often imperceptible to the human eye and difficult to faithfully reproduce. In contrast, adversarial patches have emerged as a widely adopted physical attack method against AI systems [19-hingun2023reap, 20-hu2025two, 21-wei2023unified, 22-long2025cdupatch, 23-brown2017adversarial, 24-thys2019fooling, 25-wei2024revisiting, 26-liu2018dpatch, 27-long2024papmot, zhu2024infrared, zhu2025physical, guo2025physpatch, hingun2023reap, chen2022shape, zheng2024physical, zheng2025revisiting, wei2022adversarial, wei2024infrared, wang2021dual, wang2021universal, liu2019perceptual, liu2020bias]. Early patch works focused on single-modal attacks. Thys et al. [24-thys2019fooling] demonstrate physically realizable adversarial patches that successfully mislead person detectors in the visible domain, while Zhu et al. [zhu2024infrared] propose infrared adversarial car stickers generated via a 3D mesh–shadow optimization to deceive thermal detectors. More recently, several works have extended patch attacks to visible–infrared dual-modal detectors [20-hu2025two, 22-long2025cdupatch, 21-wei2023unified]. For example, Long et al. [22-long2025cdupatch] propose CDUPatch, a color-driven cross-modal adversarial patch that achieves unified visible-infrared attacks by mapping RGB colors to corresponding thermal responses. Wei et al. [21-wei2023unified] propose a unified adversarial patch that simultaneously fools dual-modal detectors by optimizing patch shape with boundary-limited shape optimization, demonstrating high cross-modal physical-world attack success. However, existing patch attacks are always-on in everyday settings, which increases their conspicuity and complicates real-world deployment. In this paper, we co-design a thermally activated patch to mislead both visible- and infrared-spectrum detectors while remaining concealed within ordinary clothing. By leveraging temperature as a control signal, our system activates both the visible and infrared adversarial patterns simultaneously, enabling controllable, on-demand dual-modal evasion.
@@ 3.2 Structure of adversarial clothing
To enable proactive privacy protection under AI surveillance, we design an adversarial clothing featuring on-demand activation and dual-modal visible-infrared camouflage. As shown in Figure 2(a), before activation the person is detected by infrared- and visible-spectrum detectors with high confidence (0.84 and 0.85, respectively); after activation, both detectors fail to identify the person (confidence << 0.4). During this process, thermal contrast perturbs the infrared-spectrum detector, while color change deceives the visible-spectrum detector (see Supplementary Materials for the video demoxx). Figure 2(b) illustrates the layered structure of the clothing: (i) a thermochromic layer that switches from black to colorless when heated ($\sim$30 ºC), (ii) an adversarial patch layer carrying the algorithmically optimized RGB texture, (iii) a flexible silicone heating layer with precise temperature control, which activates the thermochromic layer and generates controlled infrared patterns, and (iv) a fabric substrate that provides thermal insulation for the wearer. Detailed descriptions of each layer are provided in the following subsections.
@@ 3.5 Adversarial patch layer
In the texture update phase (Figure 5b), we fix the optimized shape and learn a color texture that deceives the visible-spectrum detector. To account for real-world variations, we adopt the Expectation over Transformation [48-athalye2018synthesizing] (EOT) to simulate real-world transformations (e.g., rotation, brightness changes, blur, and scaling), thereby improving the patch’s attack effectiveness under diverse physical conditions. Then, the transformed patch is directly pasted onto RGB images to form visible-spectrum adversarial images. Finally, these patched images are fed into the visible detector, and the patch texture is updated via adversarial losses. We adopt a combination of total variation loss ($\mathcal{L}_{tv}$) and average precision loss ($\mathcal{L}_{ap}$) as the final adversarial objective, which is widely used in adversarial patch optimization [24-thys2019fooling, 22-long2025cdupatch, 38-wu2020making, 30-hu2022adversarial].
@@
Figure 6: Comparison of attack success rates (ASR) for various adversarial patch methods on infrared and visible datasets. Columns from left to right: T-SEA [36-huang2023t], AdvYOLO [24-thys2019fooling], CAP [25-wei2024revisiting], FDA [37-cheng2024full], AdvTexture [30-hu2022adversarial], AdvCloak [38-wu2020making], our RGB patch, AdvIC [31-hu2024adversarial], AdvIB [32-hu2024adversarial], AIP [33-wei2023physically], HIC-IR [34-zhu2023hiding], Bulb [35-zhu2024hiding], our infrared patch. It shows that our dual-modal patch achieves the highest attack success rate compared to other representative patch methods.
@@ 4.1 Experimental Settings
Datasets and victim models. For digital attack experiments, we evaluate our adversarial patch on person detection benchmarks using the FLIR [40-flir_dataset], INRIA [41-dalal2005histograms], PennyFudan [42-wang2007object], LLVIP [39-jia2021llvip] datasets. The FLIR and LLVIP dataset are used to assess attack performance under infrared spectrum detection, while the INRIA and PennFudan datasets are used for evaluation in the visible spectrum. We evaluate our attack on three mainstream detection paradigms: one-stage detectors (i.e., YOLOv3 [43-redmon2018yolov3], YOLOv5 [44-jocher2020yolov5]), two-stage detectors (i.e., Faster R-CNN [45-ren2015faster]) and transformer-based detectors (i.e., DETR [46-carion2020end]). To ensure accurate detection of both infrared- and visible-spectrum pedestrian targets, we use the officially pre-trained weights using the COCO [47-lin2014microsoft] dataset as the initial weights and then fine-tune these models on the training sets of INRIA and FLIR datasets. After fine-tuning, all detection models achieve over 90% mean Average Precision (mAP) on the their validation set.
@@
Figure 7: Real-world testing in diverse AI surveillance scenarios, including indoor room, shopping mall, and outdoor street. This demonstrates that our adversarial method achieves effective dual-modal evasion on both visible and infrared detectors across diverse scenarios. Refer to Supplementary Materials for video demos.
@@ 4.2 Main Results
Quantitative analysis. Figure 6 compares the performance of various adversarial patch methods on the INRIA [41-dalal2005histograms], PennyFudan [42-wang2007object], FILR [40-flir_dataset] and LLVIP [39-jia2021llvip] Datasets. As can be seen, our dual-modal patch consistently achieves the highest attack success rate (ASR) across four mainstream detectors (i.e., Yolov3, Yolov5, FastRCNN, DETR), outperforming both visible patch methods (i.e., T-SEA [36-huang2023t], AdvYOLO [24-thys2019fooling], CAP [25-wei2024revisiting], FDA [37-cheng2024full], AdvTexture [30-hu2022adversarial], AdvCloak [38-wu2020making]) and infrared patch methods (AdvIC31, AdvIB [32-hu2024adversarial], AIP [33-wei2023physically], HIC-IR [34-zhu2023hiding], Bulb [35-zhu2024hiding]). For example, on the INRIA dataset under visible-spectrum detection, our patch attains ASRs of 54.4% (YOLOv3), 49.9% (YOLOv5), 46.8% (Faster R-CNN), and 38.4% (DETR), surpassing AdvYOLO (34.6%, 31.4%, 29.7%, and 25.2%, respectively). On the LLVIP dataset under infrared detection, our patch achieves 85.4%, 88.6%, 83.3%, and 80.2%, compared to AdvIB’s 72.7%, 75.1%, 70.3%, and 65.8%. Unlike prior square-shaped patches (e.g., CAP [25-wei2024revisiting], AdvYOLO [24-thys2019fooling]), our design employs an algorithmically optimized polygonal geometry. This irregular shape enhances patch effectiveness against both visible- and infrared-spectrum detectors, allowing superior performance across diverse datasets and complex environments.